# Intrinsic Heisenberg-type lower bounds on spacelike hypersurfaces in general relativity

### v2

@@ 3.1 Universal Hardy baseline on weakly mean-convex balls
Definition 3.5 (Weakly mean-convex domain).
Let $(\Sigma,h)$ be a Riemannian manifold and $\Omega\subset\Sigma$ a bounded $C^{2}$ domain. We say that $\Omega$ is weakly mean-convex if the mean curvature $H$ of $\partial\Omega$ with respect to the outward unit normal is nonnegative in the distributional sense. Equivalently, the boundary-distance function $d(x)=\operatorname{dist}_{h}(x,\partial\Omega)$ is distributionally superharmonic on $\Omega$ (i.e., $\Delta_{h}d\geq 0$) – see [2, Thm. 1.2] and [3].
@@
On weakly mean-convex $C^{2}$ domains the boundary-distance function is superharmonic in the distributional sense, and the sharp boundary-distance Hardy inequality with optimal constant $1/4$ holds; specialized to $B_{\Sigma}(p,r)$ this gives
$\int_{B_{\Sigma}(p,r)}\|\nabla^{h}u\|^{2}\,\mathrm{d}\mu_{h}\;\geq\;\frac{1}{4}\int_{B_{\Sigma}(p,r)}\frac{|u|^{2}}{d(x)^{2}}\,\mathrm{d}\mu_{h},$ (20)
@@
$\int_{B_{\Sigma}(p,r)}\|\nabla^{h}u\|^{2}\,\mathrm{d}\mu_{h}\;\geq\;\frac{1}{4r^{2}}\int_{B_{\Sigma}(p,r)}|u|^{2}\,\mathrm{d}\mu_{h}.$ (21)
@@
Consider a slice $(\Sigma,h)$ of constant sectional curvature $\kappa>0$ (e.g. the round $S^{3}$). For geodesic balls $B_{\Sigma}(p,r)$ one has weak mean-convexity for $0<r\leq\pi/(2\sqrt{\kappa})$; hence Corollary 3.6 applies and yields the universal baseline (19). By contrast, the model-specific estimate in [10], for $\kappa>0$,
@@
is valid on the larger interval $0<r<\pi/\sqrt{\kappa}$ and decreases to 0 as $r\uparrow\pi/\sqrt{\kappa}$ (antipodal radius), where mean-convexity fails. On the overlap $0<r\leq\pi/(2\sqrt{\kappa})$ the bound (22) is strictly stronger than (19); for instance, at $r=\pi/(2\sqrt{\kappa})$ it yields $\sigma_{p}\,r\geq\pi\hbar\sqrt{3}/2\gg\hbar/2$. Thus there is no contradiction: (19) is a foliation-independent, boundary-driven floor available under mean-convexity, while (22) is a curvature-sensitive refinement that extends beyond the convexity radius.

### v4

@@ 3.1 Universal Hardy baseline on weakly mean-convex balls
Definition 3.5.
[Weakly mean-convex domain] Let $(\Sigma,h)$ be a Riemannian manifold and $\Omega\subset\Sigma$ a bounded $C^{2}$ domain. We say that $\Omega$ is weakly mean-convex if the mean curvature $H$ of $\partial\Omega$ with respect to the outward unit normal is nonnegative in the distributional sense. Equivalently, the boundary-distance function $d(x)=\operatorname{dist}_{h}(x,\partial\Omega)$ is distributionally superharmonic on $\Omega$ (i.e., $\Delta_{h}d\geq 0$) – see [15, Thm. 1.2] and [16].
@@
On weakly mean-convex $C^{2}$ domains the boundary-distance function is superharmonic in the distributional sense, and the sharp boundary-distance Hardy inequality holds: for every $u\in H^{1}_{0}(\Omega)$,
$\int_{\Omega}\|\nabla^{h}u\|_{h}^{2}\,d\mu_{h}\;\geq\;\frac{1}{4}\int_{\Omega}\frac{|u|^{2}}{d(x)^{2}}\,d\mu_{h},$
where $d(x):=\operatorname{dist}_{h}(x,\partial\Omega)$ (see [15, 16, 17]). Specialized to $\Omega=B_{\Sigma}(p,r)$ this gives
$\int_{B_{\Sigma}(p,r)}\|\nabla^{h}u\|^{2}_{h}\,\mathrm{d}\mu_{h}\;\geq\;\frac{1}{4}\int_{B_{\Sigma}(p,r)}\frac{|u|^{2}}{d(x)^{2}}\,\mathrm{d}\mu_{h},$ (27)
@@
$\int_{B_{\Sigma}(p,r)}\|\nabla^{h}u\|^{2}_{h}\,\mathrm{d}\mu_{h}\;\geq\;\frac{1}{4r^{2}}\int_{B_{\Sigma}(p,r)}|u|^{2}\,\mathrm{d}\mu_{h}.$ (28)
@@
Consider a slice $(\Sigma,h)$ of constant sectional curvature $\kappa>0$ (e.g. the round $S^{3}$). For geodesic balls $B_{\Sigma}(p,r)$ one has weak mean-convexity for $0<r\leq\pi/(2\sqrt{\kappa})$; hence Corollary 3.6 applies and yields the universal baseline (26). By contrast, the curvature-dependent estimate in [25], for $\kappa>0$,
@@
is valid on the larger interval $0<r<\pi/\sqrt{\kappa}$ and decreases to 0 as $r\to\pi/\sqrt{\kappa}$ (antipodal radius), where mean-convexity fails. On the overlap $0<r\leq\pi/(2\sqrt{\kappa})$ the bound (29) is strictly stronger than (26); for instance, at $r=\pi/(2\sqrt{\kappa})$ it yields $\sigma_{p}\,r\geq\pi\hbar\sqrt{3}/2\gg\hbar/2$. Thus there is no contradiction: (26) is a foliation-independent, boundary-driven floor available under mean-convexity, while (29) is a curvature-sensitive refinement that extends beyond the convexity radius.